OpenStudy (anonymous):

How do you change ln(1/e^4)=-4 into equivalent expression involving an expression?

OpenStudy (anonymous):

*exponent

OpenStudy (anonymous):

expand log using quotient rule for logs \[\ln \frac{1}{e ^{4}}=\ln 1 - \ln e ^{4}\]

OpenStudy (anonymous):

ln1=0

OpenStudy (anonymous):

hmmm.... doesnt' get equivalent statement... just an identity

OpenStudy (anonymous):

well i think your first response is the answer

OpenStudy (anonymous):

I think that they want to use the definition to make a statement using exponentials.

OpenStudy (anonymous):

ln e^-4 = -4 ( always true)

OpenStudy (anonymous):

But the answers look like this.

hartnn (hartnn):

ok then use this \[\log_{b}a= c \] then \(b^c=a\)

hartnn (hartnn):

and ln has base e

OpenStudy (anonymous):

yeah, that is the definition

hartnn (hartnn):

\[\log_{e}(1/e^4)=-4 \] will give \(1/e^4=e^{-4} \)

OpenStudy (anonymous):

so the exponential form should start with ... yeah that

hartnn (hartnn):

u got that @lizzzzzzzzzzz ?

OpenStudy (anonymous):

yes thank you :P

hartnn (hartnn):

welcome :)